Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2004-325_2004-331

Mission Operations Days: 2004/325 to 2004/331
November 20, 2004 00:00:00 GMT to November 26, 2004 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

AM-1 "Terra" is in Normal Mode
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
ComponentState(s)Comment
BlackbodyA On (290K); B OffNominal Science Mode
Calibration ElectronicsA On; B OffNominal Science Mode
Control ProcessorA On; B OffNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ScreenedPermanent Configuration
FDDI FormatterA On; B OffNominal Science Mode
FDDI PortA On; B OffNominal Science Mode
FIFO Memory#1 & #2 On; #3 & #4 OffNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 On, #2 OffNominal Science Mode
PV VIS FPAA On; B OffNominal Science Mode
PV NIR FPAA On; B OffNominal Science Mode
PV SMIR FPAA On; B OffNominal Science Mode
PV LWIR FPAA On; B OffNominal Science Mode
PC LWIR FPAA On; B OffNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOn (83K)Nominal Science Mode
RC SMIR FPA HeaterOffNominal Science Mode
Scan AssemblyA On; B OffNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeaterEnabledNominal Science Mode
PS2 Survival HeaterEnabledNominal Science Mode
Timing GeneratorA On; B OffNominal Science Mode
Flight SoftwareRev BD + 11 sets of patches
  • CP07 to Survival Mode - No SDD to closed command (1 patch)
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
  • CP07 to Safe Mode - No SDD to closed command (1 patch)
  • CP07 to Science Mode - Turn on FRB instead of FRA (1 patch)
  • CP07 to Outgas and Science Mode - Turn on FRB instead of FRA (1 patch)
  • FR Error Counting and Statistics (2 patches)
  • FR 1000ms time code patches (3 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
  • FR A EEPROM KME Fast 1 word patch (1 patch)
  • FR B EEPROM KME Fast 1 word patch (1 patch)
Nominal Science Mode
Inhibit Ids SetnoneNominal Science Mode
TMON 66DisabledNominal Science Mode
TMON 67DisabledNominal Science Mode
TMON 68EnabledNominal Science Mode
TMON 69DisabledNominal Science Mode

This Week’s Completed MODIS Activities:

Saturday, November 20, 2004

Extra 20 minutes of DAY rate during ground track night
2004/325 20:13:09-20:50:37 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate
2004/325 20:50:37-21:10:00 ATC MOD_OA04_NIGHT_RATE.1 OA-04: Formatter to Night Rate (10 min before groundtrack night)
2004/325 21:10:00-21:30:00 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate (night ground track)
2004/325 21:30:00-22:00:03 ATC MOD_OA04_NIGHT_RATE.1 OA-04: Formatter to Night Rate (return to normal night rate)
2004/325 22:00:03-22:41:32 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate (10 min after groundtrack day)

Sunday, November 21, 2004

None

Monday, November 22, 2004

2004/327 17:30:00-17:30:31 ATC MOD_SRCA_14_6_4_1_SETUP.1 Turns on the SRCA and sets it's lamp config. to use 10W lamp #2
2004/327 17:38:51-19:30:16 ATC MOD_SRCA_SPECIAL_10W_OPEN.2 Performs SRCA 10W continuous radiometric activity with the filter in the OPEN position
Data collect from 17:39:55-19:30:16

Tuesday, November 23, 2004

2004/328 08:00:00-08:00:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/328 08:07:26-09:05:37 ATC MOD_OA23_SRCA_FULL_SPAT_A.4 OA-23: SRCA Full Spatial Calibration - A Side
Data collection from 08:23:20-09:04:32.
S/C Night from 08:30:20-09:04:13.

Wednesday, November 24, 2004

None

Thursday, November 25, 2004

None

Friday, November 26, 2004

None

This Week’s Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Calibrations for November will include biweekly SDSM, monthly 10W Radiometric, and bimonthly SRCA Full Spatial calibrations.

Tuesday, November 30, 2004

2004/335 20:42:55-20:51:43 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/335 21:25:00-21:25:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/335 21:32:17-21:53:12 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #1
Data Collect from 21:48:11-21:53:12

Wednesday, December 1, 2004

Lunar Roll Maneuver (-8.6983 deg roll)
2004/336 20:41:29-21:17:44 ATC MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/336 21:17:44-21:24:30 ATC MOD_OA08_ROLL_MAN
2004/336 21:37:40-22:23:45 ATC MOD_OA03_FR_DAY_RATE (03:23 late)

Maneuvers:

The December Lunar Calibration Roll will be on day 2004/336 (December 1, 2004). Data collect will be from 21:17:44-21:24:30 GMT with a roll angle of –8.6983 degrees.

MODIS Anomalies:

SRCA Lamp #2 failure:

On day 2004/327 (11/22/04), Terra MODIS performed a special SRCA 10W continuous calibration with the filter in the OPEN position. This calibration was run for a duration of 1 orbit plus 10 minutes, not including lamp warmup. It was to use 10W lamp #2. However, it was noticed by MCST that this lamp shut down approximately halfway through it's scheduled on time (55 minutes, 3 seconds of the 109 minute duration). Analysis shows that there were no CP or FR event messages, and no commands to turn the lamp off. The current dropped to what would be expected if the lamp were to shut down. The SRCA stayed on until scheduled completion. Note that lamp #2 is the lamp that showed inconsistencies and was replaced by the backup lamp #4. Analysis is still being performed to determine the exact cause of this unexpected shutdown and course of action to be taken.

MODIS NAD (Nadir Aperture Door) Anomaly:

Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.

General Instrument Comments:

MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.

MODIS Telemetry Trends:

Nominal telemetry trending is ongoing.

Non-MODIS Significant Events:

Terra Solid State Recorder Anomaly

Overview:

The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA’s that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous report am1weekly_2003-305_2003-311.doc. The Terra FOT is currently evaluating plans to perform a power recycling of the SSR in an attempt to regain functionality of the “bad” PWAs. This activity will be performed at a time when it is least detrimental to critical science data.

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 237.9 of 500 hours
SRCA 10W Lamp #2: 173.3 of 500 hours
SRCA 10W Lamp #3: 182.5 of 500 hours
SRCA 10W Lamp #4: 67.2 of 500 hours

SRCA 1W Lamp #1: 567.2 of 4000 hours
SRCA 1W Lamp #2: 278.7 of 4000 hours

Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements